Lin Corporation has a single product whose selling price is $140 per unit and whose variable expense is $70 per unit. The company’s monthly fixed expense is $31,800.

Required:

1. Calculate the unit sales needed to attain a target profit of $6,700. (Do not round intermediate calculations.)

2. Calculate the dollar sales needed to attain a target profit of $9,600. (Round your intermediate calculations to the nearest whole number.)

1. Units sales to attain target profit
2. Dollar sales to attain target profit
Answer #1

Contribution margin=Sales-Variable cost  

=(140-70)=$70 per unit

Contribution margin ratio=Contribution margin/Sales

=70/140=0.5

a.Target contribution margin=Fixed cost+Target profit

=(31800+6700)=$38500

Hence target sales=Target contribution margin/contribution margin per unit

=(38500/70)

=550 units.

b.Target contribution margin=Fixed cost+Target profit

=(31800+9600)=$41400

Hence target sales=Target contribution margin/contribution margin ratio

(41400/0.5)

=$82800